# Panoramic Distortion-Aware Tokenization for Person Detection and Localization Using Transformers in Overhead Fisheye Images

### v2

@@ 4.2.2 Person detection on CEPDOF
Metrics and protocols. CEPDOF [10] provides bounding box annotations without person locations in the world coordinates. Therefore, we evaluated the detection performance using the AP. Similar to prior works [10, 48, 64], we also used precision, recall, and the F1-score. Following the procedure of [64], we trained the method of Tamura et al. [47], along with RAPiD [10], OARPD [34], Yang et al. [64], and our method, using HABBOF [22] and WEPDTOF [48] after pretraining these methods on COCO [24], although MW-R [10] was not used for the training because it is not publicly available.
Results. Our method achieved the highest mAP, AP50, and AP75 values on CEPDOF among the methods listed in Table 3. The mAP, AP50, and AP75 values achieved using our method are higher than those obtained using Yang et al.’s method [64] by values of 3.1, 7.3, and 1.6, respectively. Additionally, the F1-score of our method was greater than that of Yang et al.’s method [64] by 9.1.
@@ 4.2.3 Person detection on WEPDTOF
Metrics and protocols. We used the evaluation metrics of CEPDOF, as described in Section 4.2.2. We trained the methods of Tamura et al. [47], RAPiD [10], OARPD [34], Yang et al. [64], and our method, using HABBOF and CEPDOF after they were pretrained on COCO following [64], although MW-R was again not used for the training because it is not publicly available.
Results. Similar to the CEPDOF results, our method achieved the highest mAP, AP50, and AP75 results on WEPDTOF among the methods listed in Table 3. The mAP values achieved on CEPDOF using both the conventional methods and our method were larger than those obtained on WEPDTOF because WEPDTOF has more scenes; i.e., WEPDTOF consists of 16 videos, and CEPDOF consists of 8 videos. Comparison between the CEPDOF and WEPDTOF results shows that the methods of Tamura et al. [47], RAPiD [10], OARPD [34], and Yang et al. [64] apparently require diverse images because the network is trained on fisheye images that include rotated persons.
@@ 4.2.4 Diagnostic study
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline represents DAB-DETR [27] with the SG-Former-S [37] for the backbone when using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panoramic images were suitable for DAB-DETR because they removed the person rotations. After the panoramic conversion was adopted, the performance increased by an mAP of 9.6. We also validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided substantially improved person detection results when compared with our method without PDAT by an mAP of 4.7. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to the mAP, the other metrics were improved notably. In addition, Table 5 shows the dependence of the panoramic image width. The 3072-pixel panoramic image width is suitable for the 1024$\times$1024-pixel fisheye images because the circumference of the fisheye image is $\pi\times 1024=3217\approx 3072$. Therefore, the 3072-pixel image width without shrinkage led to the best performance among the widths of 2048, 2560, and 3072 pixels.
Table 6: Person localization results using position errors (PE) $\downarrow$ in meters on the LOAF val and test sets
Method    mPE PEn PEm PEf PEseen PEunseen
val
Seidel et al. [41] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382
Li et al. [22] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962
  Tamura et al. [47] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821
RAPiD [10] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757
OARPD [34] MTA’24 - - - - - -
Yang et al. [64] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402
Ours 0.286 0.096 0.182 0.635 0.256 0.306
test
Seidel et al. [41] VISAPP’19 1.321 0.706 1.309 3.482 1.306 1.386
Li et al. [22] AVSS’19 0.913 0.543 0.884 2.780 0.904 0.998
  Tamura et al. [47] WACV’19 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [10] CVPRW’20 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [34] MTA’24 - - - - - -
Yang et al. [64] ICCV’23 0.374 0.148 0.375 0.819 0.326 0.398
Ours 0.277 0.115 0.261 0.528 0.282 0.274
@@ 4.2.5 Person localization
Metrics and protocols. To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [64], as described in Section 2. We used the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. We also used the same subscripts used in Section 4.2.1, e.g., PEn. These estimated bounding boxes were the same as the person detection results on LOAF from Section 4.2.1.
Results. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [64] on the LOAF test set by 0.097. Similar to the person detection performance, our method realized a noteworthy 0.124 gain in terms of the PEunseen when compared with Yang et al.’s method [64]. Our method outperformed the other methods on PEn, PEm, and PEf because of the accuracy of bounding box detection.
@@ 5 Conclusion
Limitations. We assume that the fisheye images are stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using deep single image camera calibration techniques [4, 54] that estimate the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.
We have proposed a transformer-based detection and localization method for use with overhead fisheye images. Our nonuniform tokenization approach using self-similarity figures leverages the significance values to alleviate the imbalance in the significant areas of the feature maps. Extensive experiments demonstrated that our method outperforms the conventional methods substantially.

### v3

@@ 4.3.2 Person detection on CEPDOF
Table 6: Person localization results using position errors (PE) $\downarrow$ in meters on the LOAF validation and test sets
Method val test
   mPE PEn PEm PEf PEseen PEunseen    mPE PEn PEm PEf PEseen PEunseen
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382 1.321 0.706 1.309 3.482 1.306 1.386
Li et al[Li_AVSS_2019] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962 0.913 0.543 0.884 2.780 0.904 0.998
Tamura et al[Tamura_WACV_2019] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [Duan_CVPRW_2020] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [Qiao_MTA_2024] MTA’24 - - - - - - - - - - - -
Yang et al[Yang_ICCV_2023] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402 0.374 0.148 0.375 0.819 0.326 0.398
Ours 0.286 0.096 0.182 0.635 0.256 0.306 0.277 0.115 0.261 0.528 0.282 0.274
Metrics and protocols. CEPDOF [Duan_CVPRW_2020] provides bounding box annotations without person locations in the world coordinates. Therefore, we evaluated the detection performance using the AP. Similar to prior works [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023], we also used precision, recall, and the F1-score. Following the procedure of [Yang_ICCV_2023], we trained the method of Tamura et al[Tamura_WACV_2019], along with RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], Yang et al[Yang_ICCV_2023], and our method, using HABBOF [Li_AVSS_2019] and WEPDTOF [Tezcan_WACV_2022] after pretraining these methods on COCO [Lin_ECCV_2014], although MW-R [Duan_CVPRW_2020] was not used for the training because it is not publicly available.
Results. Our method achieved the highest mAP, AP50, and AP75 values on CEPDOF among the methods listed in Table 3. The mAP, AP50, and AP75 values achieved using our method are higher than those obtained using Yang et al.’s method [Yang_ICCV_2023] by values of 3.1, 7.3, and 1.6, respectively. Additionally, the F1-score of our method was greater than that of Yang et al.’s method [Yang_ICCV_2023] by 9.1.
@@ 4.3.3 Person detection on WEPDTOF
Metrics and protocols. We used the evaluation metrics of CEPDOF, as described in Section 4.3.2. We trained the methods of Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], Yang et al[Yang_ICCV_2023], and our method, using HABBOF and CEPDOF after they were pretrained on COCO following [Yang_ICCV_2023], although MW-R was again not used for the training because it is not publicly available.
Results. Similar to the CEPDOF results, our method achieved the highest mAP, AP50, and AP75 results on WEPDTOF among the methods listed in Table 3. The mAP values achieved on CEPDOF using both the conventional methods and our method were larger than those obtained on WEPDTOF because WEPDTOF has more scenes; i.e., WEPDTOF consists of 16 videos, and CEPDOF consists of 8 videos. Comparison between the CEPDOF and WEPDTOF results shows that the methods of Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], and Yang et al[Yang_ICCV_2023] apparently require diverse images because the network is trained on fisheye images that include rotated persons.
@@ 4.3.4 Diagnostic study
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline represents DAB-DETR [Liu_ICLR_2022] with the SG-Former-S [Ren_ICCV_2023] for the backbone when using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panoramic images were suitable for DAB-DETR because they removed the person rotations. After the panoramic conversion was adopted, the performance increased by an mAP of 9.6. We also validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided substantially improved person detection results when compared with our method without PDAT by an mAP of 4.7. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to the mAP, the other metrics were improved notably.
In addition, Table 5 shows the dependence of the panoramic image width. The 3072-pixel panoramic image width is suitable for the $\times1024$-pixel fisheye images because the circumference of the fisheye image is $\pi\times 1024=3217\approx 3072$. Therefore, the 3072-pixel image width without shrinkage led to the best performance among the widths of 2048, 2560, and 3072 pixels.
@@ 4.3.5 Person localization
Metrics and protocols. To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [Yang_ICCV_2023], as described in Section 4.2. We used the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. We also used the same subscripts used in Section 4.3.1, e.g., PEn. These estimated bounding boxes were the same as the person detection results on LOAF from Section 4.3.1.
Results. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [Yang_ICCV_2023] on the LOAF test set by 0.097 m. Similar to the person detection performance, our method realized a noteworthy 0.124-m gain in terms of the PEunseen when compared with Yang et al.’s method [Yang_ICCV_2023]. Our method outperformed the other methods on PEn, PEm, and PEf because of the accuracy of bounding box detection.
@@ 5 Conclusion
Limitations. We assume that the fisheye images are stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using deep single image camera calibration techniques [Dal_ECCV_2024, Wakai_ECCV_2022] that estimate the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.
We have proposed a transformer-based detection and localization method for use with overhead fisheye images. Panoramic remapping and PDAT address person rotation and small-person problems in overhead fisheye images. Our nonuniform tokenization using self-similar figures leverages the significance values to alleviate the imbalance in the significance areas of the feature maps. Experiments demonstrated that our method outperforms the prior methods.